36) if f(x)=5x-1 find f(a-2)
Please help!

Answer by giveback(3)

f(x)=5(x)-1 find f(a-2)
I guess the term (a-2) confuses you.
The trick is you just need to treat it as x and in this case x=a-2
then go back to the function f(x)=5(x)-1 and just replace x with a-2
you will have f(a-2)=5(a-2)-1=5a-10-1=5a-11
The answer: f(a-2)=5a-11